# Self-compensating Light Calorimetry with Liquid Argon Time Projection Chamber for GeV Neutrino Physics

### v1

@@ VI Discussions and outlook
In this work, we reviewed LArTPC’s capability as a dual calorimeter to estimate the energy of incident particles using the ionization charge and the scintillation light. Because of the charge recombination and the light generation mechanisms involved in the energy dissipation in LAr, the light calorimetry in LArTPC is inherently self-compensating. Good compensation of $e/h$ close to 1 can be achieved in a wide range of drift E-field from 0.2–1.8 kV/cm. Using the GeV neutrinos as a case study, we demonstrate that light calorimetry can achieve a comparable energy resolution as the more sophisticated charge imaging calorimetry, which requires pattern recognition algorithms to separate particle activities.
We made a few simplifications in this study. The LArTPC is assumed to be infinitely large so that all particle activities (except neutrinos) are contained. For finite-sized detectors, the missing energy from partially contained particles should be studied. The LArTPC is assumed to be fully active, which is not the case for detectors with multiple drift volumes. The active volume for charge and light calorimetry is assumed to be the same, but scintillation light can also be generated in LAr outside of the TPC active volume. The light yield is assumed to be uniform across the entire detector, which in practice is difficult to achieve because of the constraints of photon detector placement and achievable surface coverage. The scintillation light is assumed to be fully collected regardless of its time distribution, but the late light component could play an important role in the photon signal reconstruction and affect the calorimeter response [49]. Other contributions to energy resolution such as electronic noise, constant backgrounds, and channel-to-channel variations, are also ignored in this work. While the event-by-event fluctuation in the missing energy is often the dominating factor in the energy resolution, these other contributors could still play a large role if the detector design is not optimized for energy resolution. Therefore, the results from this work can be viewed as an optimal case for the performance of dual calorimetry in LArTPC. Dedicated simulation and analysis studies are necessary to understand all contributions to energy resolution for specific detector designs.
The results from this work can have many applications:
@@
    As a motivation and guidance to perform light calorimetry in LArTPC experiments for GeV neutrino physics. In most of the current LArTPC neutrino experiments, scintillation light is only used to provide the $t_{0}$ and trigger information. Light calorimetry to improve MeV neutrino physics has been suggested [22], but the motivation to conduct high-energy light calorimetry is lacking. This work demonstrates that in a LArTPC experiment with a capable light detection system, light calorimetry can provide an independent energy measurement with a comparable performance to the more sophisticated charge imaging calorimetry for GeV neutrinos. This will benefit numerous physics programs, such as long-baseline neutrino oscillations, atmospheric neutrinos, tau neutrino physics, and many others.
@@
    As a new approach to evaluating and mitigating the systematic uncertainties in energy measurements with LArTPCs. The systematic uncertainty associated with neutrino energy reconstruction is one of the largest uncertainties affecting the measurements of neutrino oscillation parameters and $\nu$-Ar cross sections. Two main contributors to the uncertainty in neutrino energy reconstruction are 1) mistakes in the pattern recognition algorithms when analyzing the charge-based images, and 2) dependence on the final state particle distribution provided by $\nu$-Ar interaction event generator models. The simplistic nature of the light calorimetry and the less dependence on the final state distribution (cf. Sec. IV) provides a novel way to tackle this problem.
@@
    As a foundational model to be tested. The self-compensation mechanism described in this work can be readily tested in the current and future LArTPC experiments with capable light detection systems, such as ICARUS [19], SBND [20], the DUNE LAr near detector modules [23], and the DUNE far detector prototypes [21, 24]. Care should be taken to calibrate the photon detector response and correct for effects such as the light collection non-uniformity, time integration dependence, and the selection of events.
@@
    As a design direction to improve light detection capability in future LArTPC experiments. Optimized photon detector designs, such as the APEX detector concept, are being carried out in the DUNE Phase-II program [24] for the next far detectors to improve their light detection capability such as optical coverage, photon collection efficiency, uniformity of the light response, position resolution, among others. While the original purpose for such an upgrade was to improve DUNE’s lower energy physics program such as nucleon decay, supernova neutrino detection, and solar neutrinos, the results from this work provide another strong motivation and reference on how to optimize the photon detector system design to further enhance DUNE’s capability in long-baseline neutrino oscillation and other high-energy physics programs.
@@
    As a starting point for works beyond neutrino physics. While this work focuses on the energy reconstruction of few-GeV neutrinos, it can be extended to the calorimetry of higher energy particles since the principles presented are general. The self-compensation mechanism in the LArTPC light calorimetry provides a new way to build compensating calorimeters without reducing the performance of EM responses. Coupled with the sampling calorimeter technology to contain the particle activities in a small LArTPC, scintillation light calorimetry could be considered on its own or in complementary to the well-established ionization charge calorimetry as the next-generation calorimeter for high energy collider experiments. Dedicated simulation studies are necessary to optimize its design to reach the best achievable energy resolution.
While LArTPC has been traditionally considered only as a charge-based calorimeter for GeV neutrino physics, we hope this work has proven its tremendous potential as a scintillation light calorimeter and serves as a stepping stone for future applications of light calorimetry with LArTPCs.
@@ Acknowledgements.
We thank Xin Qian and Milind Diwan for providing valuable discussions on this research. This work is supported by the US Department of Energy (DOE) Office of Science, Office of High Energy Physics under Contract No. DE-SC0012704, DE-SC0009854, BNL LDRD 23-058, and 2024 SBU-BNL Seed Grant Program.
Appendix A List of symbols
Table 1 lists the symbols used in this article, including their values, definitions, and descriptions.
Symbol Value Description
$E_{\mathrm{avail}}$ total available energy
$E_{\mathrm{dep}}$ deposited energy in LAr
$N_{e}$ (4) number of ionization electrons
$N_{\gamma}$ (4) number of scintillation photons
$N_{e}+N_{\gamma}$ total number of quanta
$\alpha$ 0.21 $N_{\gamma}/N_{e}$
$\beta$ 0.83 $1/(1+\alpha)$
$W_{\mathrm{ph}}$ 19.5 eV energy to create one quantum
$W_{\mathrm{ion}}$ 23.6 eV $W_{\mathrm{ph}}/\beta$
$R_{c}$ (6) charge recombination factor
$R_{L}$ (7) light recombination factor
$N_{e}^{r}$ $N_{e}R_{c}$
$N_{\gamma}^{r}$ $(N_{e}+N_{\gamma})R_{L}$
$E_{\mathrm{vis}}$ visible energy
$Q$ (8) visible energy in charge
$L$ (8) visible energy in light
$R_{\mathrm{cal}}$ $E_{\mathrm{vis}}/E_{\mathrm{avail}}$ calorimeter response
$e$ EM component
$h$ hadronic component
$e/h$ $R_{\mathrm{cal}}^{e}/R_{\mathrm{cal}}^{h}$
$E_{\mathrm{rec}}$ reconstructed energy
Table 1: List of symbols used in this article.
Appendix B Quanta reducing processes
There are three main processes that could reduce the total number of quanta ($N_{e}+N_{\gamma}$) in a LArTPC. However, all of them can be neglected in the scope of this work:
@@ Appendix B Quanta reducing processes
It was observed that for $\alpha$ particles, Au ions, and fission fragments at high LET, the total quanta ($N_{e}+N_{\gamma}$) collected is only about 20–70% of the ideal yield. This was explained by the biexicitonic quenching model [52], where two $\mathrm{Ar}^{*}$ atoms can combine and de-excite non-radiatively, i.e. dissipate their energy through heat. The biexicitonic quenching factor not only depends on LET, but also on the track microscopic structure such as the radial distribution of the energy deposition. While this is an important systematic effect for nuclear recoil in low energy dark matter experiments, the biexicitonic quenching was calculated to be negligible for relativistic particles and for few-MeV protons [52]. Therefore, we have neglected the scintillation quenching in our studies focusing on the calorimetry of GeV particles (and their secondaries). A second type of scintillation quenching happens when there exists parts-per-million level nitrogen contamination in LAr. In this case, the $\mathrm{Ar}_{2}^{*}$ excimer states could de-excite by collision with N2 molecules [53], thus reducing the scintillation light yield. We assume the nitrogen impurity can be well controlled and do not consider this type of scintillation quenching in this work.
@@
Appendix C Charge imaging calorimetry
@@
where $E_{\mathrm{rec}}$ is the reconstructed energy of the incident particles. In the past HEP accelerator experiments, one way to achieve this is using dual-readout calorimeters that measure both the scintillation and the Cerenkov light [55, 56]. Since most of the deposited energy in a hadronic shower is by non-relativistic particles such as low-energy protons, the Cerenkov calorimeter is less responsive to $h$. Therefore, by measuring both the scintillation and the Cerenkov light, the $f_{em}$ can be deduced event-by-event, and the $e$ and $h$’s calorimeter responses can be applied separately.
@@
Figure 10: An example 2D event display of a simulated 3 GeV $\nu_{e}$ CC interaction in LArTPC. Different colors represent the activities from different primary particles generated by the neutrino interaction.
@@
In a LArTPC, the $e$ and $h$ activities in ionization electrons can be naturally separated utilizing LArTPC’s superb 3D imaging capability with mm-scale position resolution. Figure 10 shows an example 2D event display of a simulated $\nu_{e}$-Ar CC interaction in LArTPC. Colors represent activities from different primary particles generated by the neutrino interaction. In this example, an electron, a $\pi^{0}$, and a proton are generated after the interaction, and the particles further develop into EM and hadronic showers in the detector. With dedicated LArTPC pattern recognition algorithms such as Wire-Cell [57], Pandora [58], Deep Learning [59], among others, the $e$ (e.g. the electron activities in black and the $\pi^{0}$ activities in blue) and $h$ (e.g. the proton activities in red) components can be grouped separately, and Eq. 17 can be used to reconstruct the neutrino energy. This method, also called the charge imaging calorimetry, is adopted by most of the current LArTPC experiments for energy reconstruction, which can achieve better performance than the simple charge calorimetry by adding up all charges.
@@
The charge imaging calorimetry can be further improved by correcting for the $dE/dx$ dependent recombination factor $R_{c}$ event-by-event, where the fluctuation, in particular in $h$, could be large, as illustrated in Fig. 5. The $dE/dx$ can be calculated by reconstructing all track-like objects and measuring the $dE$ for each segment of $dx$ along the reconstructed trajectories. Then, the deposited energy for track-like objects can be faithfully reconstructed using Eq. 9. However, if the track is too short, it becomes a dot-like object and the $dE/dx$ cannot be well measured since the $dx$ is difficult to reconstruct. Short tracks are commonly seen in the EM and hadronic showers that create many low-energy particles. The capability to reconstruct short tracks depends on the position resolution of the LArTPC, which typically has a pitch size of 3–5 mm and a track of a few centimeters can be well reconstructed. The separation of track-like and dot-like objects has been achieved by dedicated pattern recognition algorithms, in particular with the recent advancement in deep learning neural networks [60, 61]
Appendix D Detector smearing matrix
@@
Figure 11: The detector smearing matrices of reconstructed neutrino energy vs. true neutrino energy for electron neutrinos. The four subplots are for the four reconstruction methods Q1 (top left), Q2 (top right), Q3 (bottom left), and L1 (bottom right), respectively.

### v2

@@ VI Discussions and outlook
In this work, we have reviewed the LArTPC’s capability as a dual calorimeter to estimate particle energy from both ionization charge and scintillation light. We demonstrate that because of the recombination luminescence, the LArTPC functions as a self-compensating light calorimeter: the missing energy in the hadronic component is compensated for by the increased luminescence relative to the electromagnetic component. Using 0.5–5 GeV electron neutrino charged current interactions as a case study, we show that good compensation performance of $e/h$ from 1–1.05 can be achieved across a broad range of drift electric fields (0.2–1.8 kV/cm), with better performance for neutrino energies above 2 GeV. Under ideal conditions of uniform light collection, we show that LArTPC light calorimetry can achieve an energy resolution comparable to the charge imaging calorimetry. Challenges arising from nonuniform light collection in large LArTPCs can be mitigated with a position-dependent light yield correction derived from 3D charge signal imaging.
We made a few simplifications in this study. The LArTPC is assumed to be infinitely large so that all particle activities (except neutrinos) are contained. For finite-sized detectors, the missing energy from partially contained particles should be studied. The LArTPC is assumed to be fully active, which is not the case for detectors with multiple drift volumes. The active volume for charge and light calorimetry is assumed to be the same, but scintillation light can also be generated in LAr outside of the TPC active volume. The scintillation light is assumed to be fully collected regardless of its time distribution, but the late light component could play an important role in the photon signal reconstruction and affect the calorimeter response [56]. Other contributions to energy resolution such as electronic noise, constant backgrounds, and channel-to-channel variations, are also ignored in this work. While the event-by-event fluctuation in the missing energy is often the dominating factor in the energy resolution, these other contributors could still play a large role if the detector design is not optimized for energy resolution. Therefore, the results from this work can be viewed as an optimal case for the performance of light calorimetry in LArTPC. Dedicated simulation and analysis studies are necessary to understand all contributions to energy resolution for specific detector designs.
The results from this work can have multiple applications:
@@
    As a motivation to perform light calorimetry in LArTPC experiments for GeV neutrino physics. In most of the current LArTPC neutrino experiments, scintillation light is only used to provide the $t_{0}$ and trigger information. Light calorimetry to improve MeV neutrino physics has been suggested [22], but the motivation to conduct high-energy light calorimetry is lacking. This work demonstrates that in a LArTPC experiment with a capable light detection system, light calorimetry can provide an independent energy measurement with a comparable performance to the more sophisticated charge imaging calorimetry for GeV neutrinos. This will benefit numerous physics programs, such as long-baseline neutrino oscillations, atmospheric neutrinos, tau neutrino physics, and many others.
@@
    As a new approach to evaluating and mitigating the systematic uncertainties in neutrino energy measurements with LArTPCs. The systematic uncertainty associated with neutrino energy reconstruction is one of the largest uncertainties affecting the measurements of neutrino oscillation parameters and $\nu$-Ar cross sections. Two main contributors to the uncertainty in neutrino energy reconstruction are 1) mistakes in the pattern recognition algorithms when analyzing the charge-based images, and 2) dependence on the final state particle distribution provided by $\nu$-Ar interaction event generator models. The simplistic nature of the light calorimetry and less dependence on the final state distribution (cf. Sec. IV) provides a new way to tackle this problem.
@@
    As a foundational model to be tested. The self-compensation mechanism described in this work can be readily tested in the current and future LArTPC experiments with capable photon detection systems, such as ICARUS [19], SBND [20], the DUNE LAr near detector modules [23], and the DUNE far detector prototypes [21, 24]. Care should be taken to calibrate the photon detector response and correct for effects such as the light collection nonuniformity, time integration dependence, and the selection of events.
@@
    As a design direction to improve light detection capability in future LArTPC experiments. Optimized photon detector designs, such as the APEX detector concept, are being carried out in the DUNE Phase-II program [24] for the next far detectors to improve their light detection capability such as optical coverage, photon collection efficiency, uniformity of the light response, and position resolution, among others. While the original purpose for such an upgrade was to improve DUNE’s lower energy physics program such as nucleon decay, supernova neutrino detection, and solar neutrinos, the results from this work provide another strong motivation and reference on how to optimize the photon detector system design to further enhance DUNE’s capability in long-baseline neutrino oscillation and other high-energy physics programs.
@@
    As a starting point for works beyond neutrino physics. While this work focuses on the energy reconstruction of few-GeV neutrinos, it can be extended to the calorimetry of higher energy particles since the principles presented are general. The self-compensation mechanism in the LArTPC light calorimetry provides a new way to build compensating calorimeters without reducing the performance of EM responses. Coupled with the sampling calorimeter technology to contain the particle activities in a small LArTPC, scintillation light calorimetry could be considered on its own or as complementary to the well-established ionization charge calorimetry as the next-generation calorimeter for high energy collider experiments. Dedicated simulation studies are necessary to optimize its design to reach the best achievable energy resolution.
While LArTPC has been traditionally considered only as a charge-based calorimeter for GeV neutrino physics, we hope this work has proven its potential as a scintillation light calorimeter and serves as a stepping stone for future applications of light calorimetry with LArTPCs.
@@ Acknowledgements.
We thank Xin Qian and Milind Diwan for providing valuable discussions on this research. This work is supported by the U.S. Department of Energy (DOE) Office of Science, Office of High Energy Physics under Contracts No. DE-SC0012704, No. DE-SC0009854; BNL LDRD No. 23-058; and the 2024 SBU-BNL Seed Grant Program.
Appendix A Quanta reducing processes
There are three main processes that could reduce the total number of quanta ($N_{e}+N_{ph}$) in a LArTPC. All of them can be neglected in the scope of this work:
@@ Appendix A Quanta reducing processes
It was observed that for $\alpha$ particles, Au ions, and fission fragments at high LET, the total quanta ($N_{e}+N_{ph}$) collected is only about 20%–70% of the ideal yield. This was explained by the biexicitonic quenching model [32], where two $\mathrm{Ar}^{*}$ atoms can combine and de-excite nonradiatively, i.e. dissipate their energy through heat. The biexicitonic quenching factor not only depends on LET, but also on the track microscopic structure such as the radial distribution of the energy deposition. While this is an important systematic effect for nuclear recoils in low energy dark matter experiments, the biexicitonic quenching was calculated to be negligible for relativistic particles and for few-MeV protons [32].
In GeV $\nu$-Ar interactions, heavy particles such as $\alpha$ particles or recoiled ions (Ar, Si, Cl, etc.) may be released following a nuclear breakup or secondary nuclear scattering, with kinetic energies typically below a few MeV. Figure 12 shows the impact of the scintillation quenching from these heavy particles. The top panel shows the distribution of the energy deposited by heavy particles ($E_{\mathrm{heavy}}$) divided by the total deposited energy ($E_{\mathrm{dep}}$) from $10^{4}$ simulated 3 GeV $\nu_{e}$-Ar CC interactions. The distribution peaks near zero, indicating that heavy particle contribution is small in GeV neutrino interactions. To study the scintillation quenching impact, we applied a quenching factor (QF) of 0.7 to $\alpha$ particles [30]. For recoiled heavy ions, the QF is energy dependent. At 10–100 keV recoiled energy, the QF is between 0.25 and 0.35 [27]. For MeV recoiled energy, experimental data are less abundant, and we applied a QF of 0.25 as the maximum amount of quenching and 0.7 as the minimum amount of quenching for MeV heavy ions to study their impact. In the bottom panel, the light visible energy ($L$) distribution after applying the maximum amount of quenching is shown as the dashed blue histogram. Compared to the solid blue histogram without considering quenching, the distribution is slightly broadened in the low-energy tail. The energy resolution of the light calorimetry for 3 GeV neutrino is worsened from 5.1% to 6.1% after applying the maximum amount of quenching for heavy particles (cf. Fig. 10). If we use a QF of 0.7 for MeV heavy ions instead of 0.25, the energy resolution is only worsened to 5.4%. Since the impact on energy resolution is small ($<1$%), we have neglected the scintillation quenching in the studies presented in the main text and added the discussion here as a systematic uncertainty to consider.
Figure 12: (Top) Distribution of the energy deposited by heavy particles ($E_{\mathrm{heavy}}$), such as $\alpha$ particles and recoiled nuclei, divided by the total deposited energy ($E_{\mathrm{dep}}$) from $10^{4}$ simulated 3 GeV $\nu_{e}$-Ar CC interactions. (Bottom) The dashed blue curve shows the visible energy in light (L) after applying quenching factors for heavy particles. The other three curves are the same as in Fig. 3 and are included for comparison.
A second type of scintillation quenching happens when parts-per-million level nitrogen contamination exists in LAr. In this case, the $\mathrm{Ar}_{2}^{*}$ excimer states could de-excite by collision with N2 molecules [59], thus reducing the scintillation light yield. We assume the nitrogen impurity can be well controlled and do not consider this type of scintillation quenching in this work.
@@
Appendix B Charge imaging calorimetry
@@
where $E_{\mathrm{rec}}$ is the reconstructed energy of the incident particles. In the past HEP accelerator experiments, one way to achieve this is using dual-readout calorimeters that measure both the scintillation and the Cerenkov light [61, 62]. Since most of the deposited energy in a hadronic shower is by nonrelativistic particles such as low-energy protons, the Cerenkov calorimeter is less responsive to $h$. Therefore, by measuring both the scintillation and the Cerenkov light, the $f_{em}$ can be deduced event by event, and the $e$ and $h$’s calorimeter responses can be applied separately.
@@
Figure 13: Example 2D event display of a simulated 3 GeV $\nu_{e}$ CC interaction in LArTPC. Different colors represent the activities from different primary particles generated by the neutrino interaction.
@@
In a LArTPC, the $e$ and $h$ activities in ionization electrons can be naturally separated utilizing LArTPC’s superb 3D imaging capability with mm-scale position resolution. Figure 13 shows an example 2D event display of a simulated $\nu_{e}$-Ar CC interaction in LArTPC. Colors represent activities from different primary particles generated by the neutrino interaction. In this example, an electron, a $\pi^{0}$, and a proton are generated after the interaction, and the particles further develop into EM and hadronic showers in the detector. With dedicated LArTPC pattern recognition algorithms such as Wire-Cell [53], Pandora [54], and Deep Learning [55], among others, the $e$ (e.g. the electron activities in black and the $\pi^{0}$ activities in blue) and $h$ (e.g. the proton activities in red) components can be grouped separately, and Eq. 20 can be used to reconstruct the neutrino energy. This method, also called the charge imaging calorimetry, is adopted by most of the current LArTPC experiments for energy reconstruction, which can achieve better performance than the simple charge calorimetry by adding up all charges.
@@
The charge imaging calorimetry can be further improved by correcting for the $dE/dx$ dependent recombination factor $R_{c}$ event-by-event, where the fluctuation, in particular in $h$, could be large, as illustrated in Fig. 6. The $dE/dx$ can be calculated by reconstructing all track-like objects and measuring the $dE$ for each segment of $dx$ along the reconstructed trajectories. Then, the deposited energy for track-like objects can be faithfully reconstructed using Eq. 9. However, if the track is too short, it becomes a dot-like object and the $dE/dx$ cannot be well measured since the $dx$ is difficult to reconstruct. Short tracks are commonly seen in the EM and hadronic showers that create many low-energy particles. The capability to reconstruct short tracks depends on the position resolution of the LArTPC, which typically has a pitch size of 3–5 mm and a track of a few centimeters can be well reconstructed. The separation of track-like and dot-like objects has been achieved by dedicated pattern recognition algorithms, in particular with the recent advancement in deep learning neural networks [63, 64].
Appendix C Detector smearing matrix
@@
Figure 14: Detector smearing matrices of reconstructed neutrino energy vs. true neutrino energy for electron neutrinos. The four subplots are for the four reconstruction methods Q1 (top left), Q2 (top right), Q3 (bottom left), and L1 (bottom right), respectively.